OpenStudy (anonymous):

What is the wavelength of light that is emitted when an excited electron in the hydrogen atom falls from n = 5 to n = 4?

OpenStudy (anonymous):

You can use this handy dandy equation: \[\frac{1}{\lambda} = R_{H}(\frac{1}{n_1^{2}} - \frac{1}{n_{2}^{2}})\] \[\text{Where}\ R_H = 1.09678\times10^{-2}\text{nm}^{-1}\] Try that!